OpenStudy (iwanttogotostanford):

This one is tricky:

OpenStudy (iwanttogotostanford):

@Ineedhelplz

OpenStudy (ineedhelplz):

Monthly Rate of Decay = (Annual Rate of Decay)/12

OpenStudy (ineedhelplz):

Annual Rate of Decay = 8% Monthly Rate of Decay = (Annual Rate of Decay)/12 Monthly Rate of Decay = (8%)/12 Monthly Rate of Decay = ???

OpenStudy (daniellelovee):

0.08/12

OpenStudy (ineedhelplz):

so (8%)/12 = 0.667%

OpenStudy (daniellelovee):

yeah what he/she said^

OpenStudy (ineedhelplz):

:)

OpenStudy (iwanttogotostanford):

thank you! this one was super confusing for me and i got .667 too but i thought it was wrong lol thanks!